### v1

# Position: A taxonomy for reporting and describing AI security incidents

Abstract
AI systems are vulnerable to attacks, and corresponding AI security incidents have been described. Although a collection of safety incidents around AI will become a regulatory requirement, there is no proposal to collect AI security incidents. In this position paper, we argue that a proposal should be made, taking into account the interests and needs of different stakeholders: industry, providers, users, and researchers. We thus attempt to close this gap and propose a taxonomy alongside its requirements like machine readability and link-ability with existing databases. We aim to spark discussions and enable discussion of which information is feasible, necessary, and possible to report and share within and outside organizations using AI.
AI security, AI incidents, security incident collection
1 Introduction
Work in progress, shared for feedback purposes only. Authorlist non-final.
Artificial Intelligence (AI) is vulnerable to security threats (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). Recently, evidence appeared that such attacks occur on real-world systems (Grosse et al., 2024b; McGregor, 2021; ARVA, ). Such reports include, but are not limited to, malicious AI dependencies [^1] https://avidml.org/database/, 2023-V015, bypassed malware detectors [^2] https://avidml.org/database/, 2023-V004, failing perception modules of autonomous vehicles and reverse engineering of existing models (Grosse et al., 2024b), and search engine poisoning [^3] https://incidentdatabase.ai/cite/88.
Reporting and analyzing any incidents related to AI systems is a fundamental prerequisite for adequate risk management. For non-security incidents in context of AI systems, there are existing efforts for monitoring (OECD, ; Huntr, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b, a; Walker et al., 2024). Such databases enable analysis for deeper understanding (Kumar & Sharma, 2022). While there are taxonomies that support the reporting of cybersecurity vulnerabilities in software more broadly, such as the Common Vulnerability and Exposure (CVE) program (MITRE, e; Baker et al., 1999), these taxonomies do not apply well to AI. In contrast to shareable and protected software, AI model specifics (such as architecture, learned weights, and hyperparameters) may affect vulnerabilities (Biggio & Roli, 2018; Gu et al., 2019), but may at the same time be the intellectual property or business asset of a company (Grosse et al., 2023). Additionally, the CVE program is monitors vulnerabilities. In contrast, AI model vulnerabilities are often well-known but the details of their practical exploitation are not. AI security incident reporting requires more context for understanding AI-specific vulnerabilities and exploits that are not covered by conventional incident reporting.
Meanwhile, AI incident reporting is a developing obligation according to several regulatory frameworks. The European AI Act, as exemplary horizontal legislation, defines ’serious incidents’ in context of AI systems and requires reporting to the European AI Office and national competent authorities. At the same time, there are no corresponding templates and taxonomies for AI security incident reporting. A consensus on the necessary content, scope and granularity of suchlike frameworks has not been reached from a scientific, industry and market surveillance perspective.
Our position: We argue that a reporting taxonomy for AI security incidents is necessary and (1) should contain specific information on the AI system, the incident and its implications. (2) Its structure should be aligned with existing reporting approaches and existing AI (security) governance. (3) The content should be of sufficient granularity for industrial practitioners, researchers and market surveillance authorities to derive required action or, in some cases, even reproduce the incident. The taxonomy should further be usable, and machine readable, properties that however depend on the exact implementation. We present this draft to be further discussed by the community.
Outline. We define all relevant terminology based on existing standards or works in Section 2.1. Based on these definitions, we review generally related concepts and approaches in Section 2. We then describe a potential taxonomy for AI security incident reporting in Section 3: We first collected feedback from twelve experts for a draft reporting scheme, which was then iteratively improved. Afterwards, we describe the resulting scheme in Section 4. We then discuss limitations and implications in Section 5. For example, while our taxonomy does not cover offensive AI or recursive AI usage, we outline important implementation details. These include desired properties, such as machine readability, but also reproducibility, and the use of existing frameworks for attacks and harm classifications. To conclude, we identify open research questions like the usability of AI security reporting frameworks to ease application, and the question how to extend the framework to potentially collect information on the attacker. Before we conclude with the impact statement, we discuss alternative views in Section 6.
Related Work. To the best of our knowledge, there is no established procedure for AI security incident reporting. The closest approach to our work is by Cattell & Ghosh (2024). They propose a CVE-like framework for adjudicating AI flaws more generally (e.g., incident reporting for undesired behaviors of AI systems which is not focused on security) as well as other improvements (such as extended model cards). In this sense, Cattell & Ghosh (2024) is orthogonal to our work. In addition, MITRE announced a general AI incident sharing initiative [^4] https://ai-incidents.mitre.org concurrently to our work. While their proposal bears similarities to ours, we suggest a more detailed treatment of the intended purpose, supply chains, and existing security measures.
2 Background
In this section, we first introduce relevant definitions before we review existing works in the area of describing AI systems, AI security, as well as reporting AI and security incidents.
2.1 Definitions
As AI incident reporting is also increasingly related to regulatory requirements, we seek to define all involved terms based on existing definitions where possible. The complete list of all required definitions is in Appendix A. Here, we only revisit the three most important definitions: firstly an “AI system” (the one the incident is reported of), secondly the “CIA” security triad, and thirdly an “AI security incident” (the target of our collection). We first start with the definition of the AI system, derived from the OECD (2024a) definition.
Definition 2.1.
AI system. An AI model embedded into a larger system, together with its supporting and surrounding software and hardware. An AI system can influence physical or virtual environments. Different AI systems vary in their levels of autonomy and adaptiveness after deployment.
The incident collection we propose is limited to systems falling under this definition and covers only AI security incidents. We thus also define security as defined by ISO (2018).
Definition 2.2.
Security. The preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of a system.
We further additionally review the definition of confidentiality, integrity, and availability, better known as the CIA triad, and also well established in scientific work covering AI security (Biggio & Roli, 2018).
Definition 2.3.
Confidentiality. Property that information is not made available or disclosed to unauthorized individuals, entities, or processes (ISO, 2018).
Definition 2.4.
Integrity. Property of accuracy and completeness (ISO, 2018).
Definition 2.5.
Availability. Property of being accessible and usable on demand by an authorized entity (ISO, 2018).
Based on these security terms, we define AI security based on existing definitions by the OECD (2024b) and ISO [^5] ISO/IEC TR 24028 Information technology — Artificial intelligence — Overview of trustworthiness in artificial intelligence.
Definition 2.6.
AI security incident. An AI security incident is an event, circumstance, or series of events where the confidentiality, integrity, or availability of an AI system is actually or potentially affected and which leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labor and intellectual property rights; (d) harm to property, communities or the environment.
Our proposed taxonomy to collect AI security incidents is only intended for the above-encompassed incidents. We use these definitions to avoid, for example, confusion between security and AI security incidents (Bieringer et al., 2022). Also note that the current definition above entails potentially affected, hence vulnerabilities or red teaming results. It may be desirable to formulate this more strictly later-on.
2.2 Describing AI systems
Since the reporting of security incidents should be consistent with the overall documentation of an AI system, our proposed taxonomy builds on and expands prior work on model cards. After their initial proposal by Mitchell et al. (2019), these have been widely adopted by model providers and are used by various stakeholder groups to access information about a model and conclude corresponding action. Derivative templates have been proposed as fact sheets (Arnold et al. (2019)), datasheets for datasets (Gebru et al. (2021)), method cards (Adkins et al. (2022)), risk cards (Derczynski et al. (2023)), use case cards (Hupont et al. (2024)), and AI cards (Golpayegani et al. (2024)). While most of these approaches require the documentation of general metadata about the model (e.g. model type, model version), their content related to security is oftentimes limited and only covers implemented security measures, if at all.
2.3 AI security
Considering security is however crucial, as similar to any other (digital) system, an AI system targeted by a plethora of attacks (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). In the following, we give a very high-level overview, based on the previously discussed CIA triad, of attacks later used to sanity check the proposed reporting framework.
Poisoning attacks alter the training data, samples, or labels to decrease the overall performance of the classifier (Cinà et al., 2023; Zhang et al., 2020). They, thus, target the integrity via the performance, or availability of the AI model if performance becomes too low.
Table 1: Existing databases and collections of AI (security) incidents. We list the kind of support party (Government, Industry, Non-Profit, and Academia) and incident numbers as of June 14th, 2024. We list the number of entries in each database and whether incidents cover the topics of ethics, security, AI security, or offensive use of AI within security. When then review which databases contain reproducible incidents. We then revisit the sources of the incidents: research papers, media, anonymous reports, or platforms where registered users can submit reports. We finally report whether the databases are machine-readable and linked with the CVE databases. The AVID database is linked to NVD, MITRE, and AIID.
Incident Type Sources
Database # entries
Ethical
Security
AI Security
Offensive
AI usage
Reproducible
incidents
Research
paper
Media
Anonymous
Submissions
Machine
readable
Direct
integration
Linked with/to
OECD AIM G $>$1,000 CSV
Huntr I $>$1,000 CVE
AIID (McGregor, 2021) NP 691 CSV MongoDB
Deepfake DB (Walker et al., 2024) A 200 CSV
AVID (ARVA, ) NP 48 JSON Python
LVEs A 36 JSON Python
Grosse et al. (2024b) A 32
MITRE case studies (MITRE, b) G 22 JSON Python CVE
AIRD (MITRE, a) I 4 CSV Python CVE
Evasion attacks tamper with the test input to change the output of the model (Biggio & Roli, 2018; Dalvi et al., 2004; Boloor et al., 2020; Abdelnabi et al., 2023), thus harming the model’s integrity.
Model stealing copies the model without the owner’s consent (Oliynyk et al., 2023; Chen et al., 2021; Carlini et al., ), thus targeting confidentiality.
Membership inference infers information about the training data based on the model’s replies to queries (Shokri et al., 2017; Pan et al., 2019; Fu et al., 2024). This attack thus also targets confidentiality.
Many more attacks, including backdoors (Cinà et al., 2023; Yu et al., 2022; Yang et al., 2024), weight perturbations (Neti et al., 1992; Breier et al., 2018), adversarial initialization (Grosse et al., 2020; Liu et al., 2019), and sponge attacks (Shumailov et al., 2020) have been discussed. The exact nature and scope of describing and tagging a specific attack in a practical context is a research question itself and thus left for future work. There are, however, established frameworks like the MITRE ATLAS [^6] https://atlas.mitre.org/matrices/ATLAS addressing this challenge in practice.
2.4 AI incidents
Although there are rather few reports of real-world AI security incidents as discussed in the introduction (Grosse et al., 2024b; McGregor, 2021; ARVA, ), there are many reports of general incidents in the context of AI systems. We summarize related databases in Table 1. Most databases cover not only AI security incidents but instead a range of incidents, ranging from unethical applications of AI to security, and offensive use of AI (OECD, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b). Very few databases are more specific and focus on non-AI security (Huntr, ) or offensive AI usage (Walker et al., 2024). In addition, most databases rely on publicly available information, like research papers and media reports (OECD, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b). Far fewer rely (also) on direct submissions (Huntr, ; LVEs, ; MITRE, a) in the spirit of bug bounties. Even fewer use anonymous reports (Grosse et al., 2024b). Each format constrains the kind of information that is available per database entry. While some databases, such as the AI Incident Database (AIID) (McGregor, 2021), apply taxonomies to incidents, they do not require the information needed for an in-depth security analysis that allows to determine if other AI systems are vulnerable or not. Opposed to this, MITRE recently released an AI incident sharing initiative [^7] https://ai-incidents.mitre.org collecting more information.
Still, these databases provide guidelines on collected information and emphasize other relevant features. These include machine readability (seven of nine databases allow direct export of data), links to other vulnerability databases like CVE (four of nine), and lastly direct integration via for example APIs with programming languages for data analysis (five of nine). The latter is especially relevant to research, where corresponding frameworks in non-AI security have enabled more detailed analysis of the threat surface (Neuhaus & Zimmermann, 2010; Li et al., 2017).
2.5 Existing infrastructure for security incidents
Within security, frameworks to collect vulnerabilities exists. The main reporting infrastructure for security is the CVE program (Baker et al., 1999). Similar taxonomies have evolved over time to accommodate emerging new methods and tools (MITRE, f). For example, in the case of CVE (Baker et al., 1999), a related method, the Common Vulnerability Scoring System (CVSS) (NIST, a; Scarfone & Mell, 2009), was developed to assess the severity of vulnerabilities. As a result, the National Vulnerability Database (NVD) (NIST, b) now contains CVSS scores for all published CVE records. Additionally, alongside the descriptions of vulnerabilities and exposures, the need to enumerate attacks was identified. The MITRE ATT&CK framework (MITRE, c, d), which serves as a taxonomy of attack tactics and techniques, was created to address this gap.
Previous works outline that security and AI security vulnerabilities are fundamentally different (NIST, 2023; Cattell & Ghosh, 2024). We agree with this position, yet acknowledge that CVEs can be relevant for AI models [^8] For an example, consider CVE-2023-25661.. CVEs are however orthogonal to our reporting in that they cover model independent vulnerabilities in for example code, while our taxonomy covers model specific vulnerabilities (which may extend to several models, but are not fully model agnostic).
3 Approach
To derive a taxonomy for reporting and describing AI security incidents, we followed a three step approach. Firstly, we reached out to twelve industry experts (working in AI red teaming, security industry, or AI governance), inquiring for their wish-list for a possible reporting scheme. We then compared the stated preferences with each other as well as existing incident collections or approaches similar to model cards. The derived, rudimentary scheme was then refined by iterative discussion rounds among four of the authors. When both draft and definitions had sufficiently converged, three of the authors used the tentative structure to tag incidents from existing databases. All appearing flaws were consecutively fixed, and the proposal was shared with the initial and additional experts for a broader, second round of feedback.
4 AI security incident taxonomy
Our taxonomy consists of five parts, where the last contains only an open text field for further comments and feedback, and is omitted from the following presentation. The parts are, (1) the incident’s meta data, (2) information about the targeted AI system, (3) specifics of the security incident, and lastly (4) implications of the incident. We now revisit the individual questions and references in each part. An example of the final questionnaire is depicted in Table 2-4 in the appendix.
4.1 Meta data (1.1-1.4)
The metadata is a short block containing the incident’s high-level information, and is, under variations, part of many databases (Huntr, ; McGregor, 2021; LVEs, ; ARVA, ; OECD, ). This information forms the first section with elements 1.1-1.4.
We opted to assign an incident ID, a name or ID of the incident reporter, their organization, and the alleged date (or interval, or unknown) of the incident. It is important to remark that an AI security incident may be reported anonymously. Although, as stated above, any question of our proposal can be dropped, we want to emphasize that the reporter and their organization may be pseudonymized or kept fully anonymous, depending on the specific usage of the taxonomy.
4.2 Targeted AI system
To understand an AI security incident, it is crucial to describe the AI system that formed part of the incident. Hence, in the second section of our taxonomy, we first collect information about the model, its life cycle stage and use, its origin, and application-specific information. Ideally, this information is directly imported from a model card or similar documentation to reduce required work.
Model information (2.1-2.3). We first collect basic information about the model. This includes the model type, for example, whether a model is generative or predictive; a specific reply here could be a support vector machine. Such information was suggested by half of our experts and is used in several model card approaches (Mitchell et al., 2019; Arnold et al., 2019; Adkins et al., 2022). We then collect more specific information like the exact version of the model, information also encompassed in many model cards (Mitchell et al., 2019; Adkins et al., 2022) and requested by a fourth of our experts. Lastly, if possible, the model should be specified as detailed as possible (e.g., model weights), to enhance reproducibility, as suggested by another quarter of our experts.
Life-cycle stage and intended use (2.4-2.7). In addition to the model within the AI system, understanding the system context is crucial. To this end, we also collect lifecycle information as specified by ISO [^9] ISO/IEC 22989 Information technology — Artificial intelligence — Artificial intelligence concepts and terminology, https://www.iso.org/standard/74296.html, as suggested by two of our experts. As suggested by the CSET AI Harm Framework (Hoffman & Frase, 2023) annotation guide, we also collect the domain where the AI system is deployed (Hoffmann et al., ). Intuitively, a deployment in physical domains has different implications than in the digital space. In addition, we keep track of the intended use of the AI system, loosely following the EU AI Act (Council of European Union, 2024), as done in many model card approaches (Mitchell et al., 2019; Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). To take into account specific exploits, we also advocate denoting a usage during the incident if different from the intended use.
Model supply chain (2.8-2.10). Furthermore, it is crucial to understand possible (external) influences on the AI system stemming from the supply chain, as also suggested by two of our experts and in several AI model card approaches (Mitchell et al., 2019; Gebru et al., 2021; Adkins et al., 2022). Following scientific literature of AI security, we opted to divide the supply chain into three major areas: data (Mir, 2024), the model weights (Gu et al., 2019), and the code (Grosse et al., 2020; Shumailov et al., 2021). Any of these may be reused and thus allow the import of security vulnerabilities.
Application (2.11). Instances of AI security incidents may need specific details to be reproducible or understandable. As an example, we put forward also collecting the system prompt in the case where a large language model (LLM) is used. Such information was suggested by one expert, is used in LLM-related databases (LVEs, ), and is used in model cards (Derczynski et al., 2023). Such information may be determined for individual usage, and may thus be adapted freely.
4.3 Incident information
Having characterized the AI system, we are ready to collect information about the security incident with the goal of understanding what (may have) happened. To this end, we first collect general information about the incident, then report relevant affected components and finally understand which defensive measures were in place.
General information (3.1-3.4) First, the materialization status is collected, as suggested by two experts and as done in one database (ARVA, ). This information distinguishes a vulnerability found during testing (e.g., red-teaming) from an exploit of a deployed system. We further put forward attempting to find the root cause of the incident and report it. To ensure that an incident is indeed an AI security incident as defined in Definition 2.6, the attacker’s malicious intent should recorded. This was suggested by half of our experts, forms part of several databases (ARVA, ; OECD, ; Walker et al., 2024), and was suggested in prior work (Grosse et al., 2024b). Finally, we recommend recording the attack type. For coding purpose and in this paper, we rely on the categorization by Grosse et al. (2024b). Any other framework, however, may also serve this purpose.
Relevant components (3.5-3.7) An AI security incident is further strongly categorized by the resources used by the attacker to carry out the attack (Apruzzese et al., 2023). We thus recommend identifying and reporting the target asset, e.g. whether the model or the entire system was targeted, pointed out in prior work (Bieringer et al., 2022). In addition, we propose to keep track of hardware in relation to the incident, as suggested by one expert. Finally, access rights to all components of the AI (data, model, queries, etc) should be reported to identify how the attacker possibly entered the system. The latter was suggested by one expert and studied in prior work (Grosse et al., 2024a, 2023).
Defensive measures (3.8 and 3.9) The access patterns also give an understanding how which access, in theory, should not have been possible. Along these lines, there is value in understanding what measures were in place before the incident took place, but did not prevent it. To this end, we advocate reporting existing security measures, as suggested by three experts and the literature in model cards (Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). This is in particular relevant for large language models, where implemented guardrails should be reported (Rebedea et al., 2023). Along the same lines, we suggest also collecting fail-safe measures.
4.4 Incident implications
To understand the severity of an AI security incident, it is crucial to monitor the damage it causes. Monitoring such effects is part of this brief section with elements 4.1 to 4.4. The information in this section may be difficult to derive depending on the role of the reporter—a possible solution for this problem is to rely on a committee or other help.
As the first element, we propose to monitor whether the incident led to non-compliance with existing regulations, a factor brought up by industry participants in security contexts (Bieringer et al., 2022; Grosse et al., 2023). Secondly, any harm that occurred should be well-documented, as suggested by the CSET AI Harm Taxonomy (Hoffman & Frase, 2023) and applied within the AI Incident Database (McGregor, 2021). While it is recognized that determining (or defining) the extent of harm of an AI incident can be inherently uncertain (Paeth et al., 2024), we advocate that it is still worth attempting to record. A different level of abstraction is to use the CIA definition from Section 2.1 to characterize the impact of the incident.
Finally, and beyond the immediate incident, the goal of our taxonomy is to understand vulnerabilities beyond an individual system. We thus advocate explicitly reporting an assessment of whether the incident could affect other AI systems. This was also suggested by three of our experts and is implemented in the Huntr database already (Huntr, ).
4.5 Appendix
As a last part, we propose a field for links to connect the reported incident to related incidents from either a different database or a similar incident from the same reporting scheme, via either name or an URL. Finally, we suggest adding a comment field. This allows adding relevant information missing in the previous parts. Such a field avoids loss of information while seeking to keep the reporting taxonomy directed and usable.
5 Limitations, implementation, Implications and future research
Having discussed the proposed taxonomy, we now revisit known limitations, how the taxonomy should be implemented to be of most use, and finally implications and open research questions.
Limitations. As stated in the introduction, this taxonomy is a basis for a discussion within the community, not a final version. As such, it has several limitations. Firstly, it does not cover offensive AI, as we consider this usage orthogonal to the security incidents we aim to capture. Secondly, the system currently does not handle AI systems with multiple AI models or components, or recursive usage of AI. While an extension covering several AI systems is possible by filling questions $2.x$ for all covered AI models, there may be better responses to this shortcoming. In general, the taxonomy will have to be adapted for new use cases and trends in AI. Furthermore, security incidents are very application specific [^10] As an example, consider AlphaFold’s prohibited use policy https://github.com/google-deepmind/alphafold3/blob/main/WEIGHTS_PROHIBITED_USE_POLICY.md, and it may be necessary to adapt the taxonomy for specific applications. Furthermore, our current definition of security also covers incidents with “potential” effects. While this enables to cover red-teaming results or vulnerabilities with unknown statuses of exploitation, in the long run it may be more beneficial to align the definition of security definitions including only incidents with identified harm.
Implementation. Many properties desired by the community, such as ease-of-use, machine readability (Sect. 2.4), and reproducibility (Sect. 4.2) depend on the practical implementation of our taxonomy. For example, the taxonomy is best to use when frameworks for attack and harm categorization are well known to the reporter and downstream audience. We thus suggested that external frameworks in certain corresponding questions (for example, attacks and harm questions), can be utilized (Sect. 4.3). Along these lines, machine readability (Sect. 2.4) depends strongly on whether answers are selected from finite lists (such as multiple-choice or drop-down responses) as opposed to open text fields, or defining the taxonomy keys and values in a more generalizable schema (such as XML, RDFa, or other standards). While we most likely cannot eliminate all open text fields, we strongly suggest minimizing them as much as possible in order to allow ease of analysis. Lastly, it would be good to allow the reproducibility of an incident (Sect. 4.2). However, reproducibility directly contradicts the intellectual property of a model release. In this context, it would be possible to reduce the reporting scheme to contain only information supporting understanding of the incident, and not release model weights, for example. Otherwise, a trusted third party could act as an intermediary.
Implications and future work. We have outlined that the exact reporting taxonomy and implementation strongly influence the overall usefulness and applicability of an AI incident reporting tool. Understanding the requirements of new emerging trends within AI (security), new applications, and, machine readability of AI incident data are open research questions. This extends to the tension between the protection of proprietary AI model information and the learning and generalizability of AI security incidents. The draft we present here is intended as a scaffold to determine, discuss, and apply such changes. Orthogonally, but analogously to other areas of incident reporting, usability studies could help to understand and advance both the usefulness of the collected reports (Marcilly et al., 2019; Mazaheri et al., 2015; Wiele & Rantanen, 2015). Another aspect is to report the identity of the attacker. While similar frameworks such as CVE neither cover the identification or the collection of information on the attacker, in the long run approaches obtaining such information would greatly benefit both research and AI users.
6 Alternative views
The position and suggested taxonomy in this paper is not the only way to approach AI security incident reporting. While the exact taxonomy, as stated, is open for debate, there are more extreme positions to be taken on this topic.
Position—No reporting is needed. One may argue that due to the absence of many AI security incidents, AI security reporting is not (yet) needed. In contrast, we believe that based on existing reports of AI security incidents (Grosse et al., 2024b; McGregor, 2021; ARVA, ) and the idea of proactive security (Biggio & Roli, 2018), incident collection should rather start too early than too late.
Position—CVE reporting is sufficient. One approach is integrating AI security reports into the existing CVE program. While we have highlighted several reasons against this approach (Sect. 2.5), it could instead be argued that the CVE scheme could be altered to better fit AI security incidents. We have reasoned that CVE is too general and does not consider that data on deployed AI models are often withheld by companies. However, regulatory frameworks might require model providers and deployers to make the specific information related to AI security incidents transparent.
Position—Extending general AI incident reporting. A different approach to the one described above could be using the existing infrastructure (Sect. 2.4) for non-security AI incidents and extending it to security. While we argue that the frameworks are sufficiently different and security incidents merit specific attention, existing infrastructure and AI incident communities could be extended towards security.
Position—A smaller, less bureaucratic approach is needed. There may be justified criticism that the 26 questions (Sect. 4) are too many and a bureaucratic burden for the reporters. While we have tried to argue in favor of reusing existing infrastructure such as model cards (Sect. 4), we would like to emphasize that the here listed elements are not a minimum, but an ideal suggestion and wish list of several parties. We are open to discussing which of the 26 questions could be a minimal set to still fulfill the purpose of different stakeholders.
Position–Taxonomies evolve over time. As reasoned above, security reporting infrastructures have been changing over time to accommodate emerging new methods and tools (Sect. 2.5). It is thus expected that a taxonomy of AI security incidents will evolve in a similar fashion over time, especially our proposed taxonomy. Yet, we argue that this draft is valid as a base of discussion and that eventually, a framework is needed to standardize communication between reporting entities and entities receiving reports as required legally.
Impact statement
The current annual damage of cyber crime, when quantified in money, is 8 trillion USD in 2023 alone. [^11] https://www.statista.com/forecasts/1280009/cost-cybercrime-worldwide Understanding and combating such crime is thus of uttermost importance. While virtually all of this number is currently caused by “classical” security issues (unrelated to AI), the deployment of AI systems across sectors is dramatically increasing and AI is likely to be implicated in far more cybercrime losses; AI security vulnerabilities are rather well established (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). At the same time, little is known about the security incidents that occur in the wild on deployed systems (Apruzzese et al., 2023; Grosse et al., 2024b). In aviation, the rigorous study of incidents has lead to a decrease in casualties over time (Li, 2019). Analogously, knowledge about practical AI security incidents could fuel AI security research leading to research in more practical threat models. We hope with our paper to make a significant step towards a taxonomy for AI security incidents that benefits all stakeholders—the developers and deployers affected, but also research, industry, and governmental entities who seek to analyze and learn from these incidents.
Acknowledgments The authors would like to thank Jochen Stängler (German Federal Office for Information Security, BSI) who has given extensive feedback on both the definitions and the draft of the taxonomy.
Appendix A Definitions
In the main paper, we list the most important definitions. To define all aspects of our taxonomy well, we introduce all required definitions and their sources in the following section.
Definition A.1.
Vulnerability. A weakness of an asset or a measure to reduce risk that can be exploited as a potential cause of an unwanted incident, which can result in harm to a system or organization, from ISO/IEC 27000:2018 (ISO, 2018).
Definition A.2.
Attack. Attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of an asset, from ISO/IEC 27000:2018 (ISO, 2018).
Definition A.3.
AI system security. preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of an AI system, derived from security definition (Def. 2.2).
Definition A.4.
AI security. preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of an AI model, derived from Security definition (Def. 2.2).
Definition A.5.
Incident. An incident is an event, circumstance or series of events where the development, use or malfunction of one or more systems directly or indirectly leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, from the OECD [^12] https://www.oecd-ilibrary.org/docserver/d1a8d965-en.pdf?expires=1718528495&id=id&accname=guest&checksum=8938FF35EAD65EAADD585DCF1FE65791.
Definition A.6.
AI Incident. An AI incident is an event, circumstance or series of events where the development, use or malfunction of one or more AI systems directly or indirectly leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, from the OECD [^13] https://www.oecd-ilibrary.org/docserver/d1a8d965-en.pdf?expires=1718528495&id=id&accname=guest&checksum=8938FF35EAD65EAADD585DCF1FE65791
Definition A.7.
Non-security incident. Any incident that was not caused by an attack and harmed the security of a system, based on definitions for security (Def. 2.2) and attack (Def. A.2).
Definition A.8.
Non AI-security security incident. A non AI security incident is an event, circumstance or series of events where the confidentiality, integrity, or availability of an AI system, but not the ML model, is actually or potentially affected and which leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, derived from the definitions of AI security, security, and attack (Def. A.2).
Definition A.9.
Offensive use of AI. The use or abuse of ML to accomplish a malicious task, as defined by Mirsky et al. (2023).
Definition A.10.
AI model. An AI model is a machine-based system that, for explicit or implicit objectives, infers from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions, as part 1 of the OECD definition [^14] https://oecd.ai/en/wonk/ai-system-definition-update.
Definition A.11.
Data. Re-interpretable representation of information in a formalized manner suitable for communication, interpretation or processing, from ISO/IEC TR 24028:2020.
Definition A.12.
Infrastructure. Hardware and software environment to support computer system and software design, development, and modification, from ISO/IEC TR 17903:2024.
Definition A.13.
Safety. Freedom from risk which is not tolerable, from ISO/IEC TR 5469:2024.
Table 2: Reporting scheme with description for annotators and examples, first third (Part 1.1 - Part 2.11).
Short name Long name or question Description for annotators Example
Meta Information 1.1 Incident ID Incident ID The unique identifier of the incident. ID_1716
1.2 Incident Reporter Name of incident reporter Name or also username from within recording. John Doe / AIIncidentreporter001
1.3 Reporting Organization Name of reporting organization The legal name of the reporting organization. Acme Inc.
1.4 Incident Date Date the incident ocurred Date, Time interval, if unkown estimate interval 13/12/2024
AI System Information 2.1 Model type Specify the ML model type. Generative or predictive AI? prediction, generative, but also more concrete, Transformer, DNN, ConvNet, SVM etc Support Vector Machine
2.2 Model version High level identifier, as noted in model card. This identifier should allow for correct identification of the specific model version. Llama3.2:1B
2.3 Model specification Provide as precise information on the model as possible. Information could contain URLs to or files with weights, full model files or model cards. ’URL’
2.4 Lifecycle stage Define the stage of the AI lifecycle which the model was in when the incident occurred. Stages of the AI lifecycle are: Inception, Design & Development, Verification & Validation, Deployment, Operation and Monitoring, Re-evaluate, Retirement Deployment
2.5 Domain Did the incident occurr in a physical environment? Did the incident affect an AI that is interacting with physical objects or otherwise physically with humans? Yes, pedestrian not detected.
2.6 Intended use Described the intended use of the AI system. Intended purpose means the use for which an AI system is intended by the provider, including the specific context and conditions of use, as specified in the information supplied by the provider in the instructions for use, promotional or sales materials and statements, as well as in the technical documentation. Object detector to tag cats on image, Chatbot to reply to questions about air plane fares, etc
2.7 Actual use during incident, if differing from 2.6 Describe the use of the AI system during the incident, if differeing from intended use (2.6) This may mean the usage for which an AI system was not intended by the provider, including the specific context and conditions of the actual use. Used to tag foxes instead of cats, used to get legal advice model was not trained for.
2.8 Model supply chain - Data Was publicly available data used? If yes, from where, was data extended or adapted and if yes how? Specify dataset with name and URL, as well as any modifications that the data underwent before being used. For example: Model used MNIST data from ’URL’, where we added 1000 images that were handlabelled by our team.
2.9 Model Supply Chain - Model weights Are the model weights derived from an existing model, and if yes, which one and how were they adapted? If the original design or weights have a name, provide this name or possibly a URL. Further describe all steps the model undewerwent to adapt it to the current status. For example: Based on Llama3.2:1B, but finetuned on the above described dataset for n iterations using optimzier o and loss l.
2.10 Model Supply Chain - Code What public code was used to generate and run the model? Provide libraries used with their version name, and also ennumnerate github repositories with link and access data or other re-used code from the project. For example: Pytorch version 2.4.1. and Code from github repo ’this url’ was used, forked January 1st 2024.
2.11 System prompt If the system is a generative AI system, provide the system prompt. If known, report the entire system, prompt as used in the setting where the incident occurred. For example: ’please be polite to the user and don’t hallucinate. The user writes: ’
Table 3: Reporting scheme with description for annotators and examples, second third (Part 3.1 - Part 3.9).
Short name Long name or question Description for annotators Example
Incident specification 3.1 Materialization status Has the incident been reported prior or after its materialization? Was the attack found during red-teaming or scientific analysis, or as part of an exploited breach? Red teaming
3.2 Root cause Describe the root cause of the incident. What is the underlying reason, design choice, or property of the system that enables the attack? For example: ’Attacker relied on infinity norm attack the model wasn’t tested for’, or ’API allowed infinite amount of queries, hence enabling crating a copycat training dataset’
3.3 Malicious Intent Is there evidence that an entity or a person purposely caused the incident by targeting the AI systems security? Is the incident an example of an attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of the AI system? For example: ’Yes, the attacker altered the sample until it passed the AI security check’ or ’No, the user did not know that the system was not trained on this specific subset of the training data (Disease B). Hence, the system provided a wrong output.’
3.4 Attack type Specify the concrete type of attack that has been carried out. Any attack framework, tagging based on the attacks in Sect. 2.3 Poisoning
3.5 Target asset Did the attack target the model explicitly and the model only? Based on the definition in Sect. 2.1 Yes, model was targeted.
3.6 Hardware What hardware was used, and could it have affected the incident? Please describe the specific hardware used on the device or server the application ran, and whether the setting implies that the hardware played a role in the incident. Mac M1, could not have affected the incident OR could have affected the incident via a glitch.
3.7 Access Which access was given to whom for model, inputs, training, test or evaluation data? Specify which levels of access were given to which users regarding each component. Training and evaluation data as well as model was only available to internal model team. The model could be queried via API by the client, but access control prevented any other access by 3rd parties to the model.
3.8 Security measures What security measures where used? Where they circumvented and if yes how? Give a list of specific security measures that were used including their configurations. Adversarial training using FGSM with epsilon 0.1, 0.2 and 0.3 was used on randomly selected 10% of the dataset were used. Defense was not circumvented as attack was based on L2 norm, not $L_{\infty}$ norm.
3.9 Fail safe measures Where there any fail-safe measures implemented? Describe any fail-safe measures, such as rebooting the system, not outputting a classification, etc. System fell back on human intervention to prevent further harm.
Table 4: Reporting scheme with description for annotators and examples, last third (Part 4.1 - Part 5.1).
Short name Long name or question Description for annotators Example
Incident Implications 4.1 Regulatory damage List specific regulatory requirements that are violated due to the incident. Please specify any concrete legislation, standard or policy that has been violated by the incidents’ occurrence. European AI Act
4.2 Resulting harm What harm did the incident cause to any participating stakeholder? Did the incident lead to any of the following types of harm? Affecting physical health, inducing financial loss, damaging a physical object, affecting intangible property, affecting infrastructure, affecting natural environment, violation of human rights, civil liberties, civil rights or democratic norms. Harm to physical property of all users, mainly food but also some devices due to malfunction of the fridge AI.
4.3 Security Impact Describe how the incident affected the confidentiality, integrity, availability of the AI system. Describe whether the incident caused any leakage of confidential data, lead the system to output a wrong result or to exhibit a bias (integrity), or lead to the system being down or otherwise unusable (availability). During the incident, samples were manipulated to be classified as vehicles and thus harmed integrity.
4.4 Generalization Does the incident potentially affect other AI systems? Incident descriptions should allow to determine for a third model if a vulnerability in that model is likely. Other AI systems might also be affected if relying on the same dataset ”xyz”.
Appendix 5.1 Links to other incidents Links to similar / same cases in other /same database(s) In case you are aware of a incidents that has been reported in another database or that is similar to this incident in this database For example, this incident has had similar consequences like AIID incidents 38 or simply ’URL’
5.2 Comments Additional comments. Anything you’d like to add which has not been covered in the previous fields. For example, the attack does not work on model ABC1.2 and ABC1.4, but only on ABC1.3