### v2

# Position: A taxonomy for reporting and describing AI security incidents

Abstract
As AI usage becomes more ubiquitous, AI incident reporting is both practiced increasingly in industry and mandated by regulatory requirements. At the same time, it is established that AI systems are exploited in practice by a growing number of security threats. Yet, organizations and practitioners lack necessary guidance in describing AI security incidents. In this position paper, we argue that specific taxonomies are required to describe and report security incidents of AI systems. In other words, existing frameworks for either non-AI security or generic AI safety incident reporting are insufficient to capture the specific properties of AI security. To demonstrate our position, we offer an AI security incident taxonomy and highlight relevant properties, such as machine readability and integration with existing frameworks. We have derived this proposal from interviews with experts, aiming for standardized reporting of AI security incidents, which meets the requirements of affected stakeholder groups. We hope that this taxonomy sparks discussions and eventually allows the sharing of AI security incidents across organizations, enabling more secure AI.
AI security, AI incidents, security incident collection
1 Introduction
Reporting and analyzing any incidents related to Artificial Intelligence (AI) systems is a fundamental prerequisite for adequate risk management. For non-security incidents in the context of AI systems, there are existing efforts for monitoring incidents (OECD, b; Huntr, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a, ; Walker et al., 2024). In other areas like aviation, the rigorous study of incidents has lead to a decrease in casualties over time (Li, 2019). Consequently, AI incident reporting is becoming a relevant obligation according to regulatory frameworks. The European AI Act, as exemplary horizontal legislation, requires reporting of serious incidents to the European AI Office and national competent authorities. Yet, this entails a broad range of incidents: general incidents as above, but also incidents where an attacker manipulates an AI system.
There is a plethora of scientific literature on how to attack or manipulate AI (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). Recently, evidence appeared that such attacks occur on real-world systems (Grosse et al., 2024b; McGregor, 2021; ARVA, ). Such reports include, but are not limited to, malicious AI dependencies [^1] https://avidml.org/database/, 2023-V015, bypassed malware detectors [^2] https://avidml.org/database/, 2023-V004, reverse engineering of existing models (Grosse et al., 2024b), and search engine poisoning [^3] https://incidentdatabase.ai/cite/88. Research confirms (Carlini et al., 2024a) such threats for widely deployed AI systems like large language models (LLMs) (Baier, 2024), and text-based attacks can be mined from the internet and thus manifest there as well (Shen et al., 2024). Yet, AI security incidents, as of now, seem not to occur widely (Grosse et al., 2024b). It would thus be an ideal moment to start collecting AI security incidents for a proactive, not reactive, approach.
Intuitively, an attacker warrants different information to be collected than other incidents. Examples for incident collection from non-AI security include the Common Vulnerabilities and Exposures (CVE) program (MITRE, d; Baker et al., 1999). CVE is an established industry standard, and similar to the aviation case above, the collected data enable analysis for deeper understanding (Kumar & Sharma, 2022), like for example trends in of frequencies of exploits. However, existing security taxonomies do not apply well to AI: In contrast to shareable and protected software, AI model specifics (such as architecture, learned weights, and hyperparameters) may affect vulnerabilities (Biggio & Roli, 2018; Gu et al., 2019). Model specifics may also be the intellectual property or business asset of the affected company (Grosse et al., 2023). Furthermore, taxonomies like the CVE program catalog vulnerabilities found in the wild. In contrast, AI model vulnerabilities are often well-known with unknown practical exploit status. Understanding the exact exploit would be more fruitful. AI security incident reporting should thus require more context for understanding AI-specific vulnerabilities and exploits not covered by conventional incident reporting. We thus need a broad consensus on the necessary content, scope and granularity of AI security incident reporting that has not yet been reached from scientific, industry and market surveillance perspective.
Our position: A reporting taxonomy for AI security incidents is necessary and (1) should contain specific information on the AI system, the incident and its implications. (2) Its structure should be aligned with existing reporting approaches and regulatory disclosure obligations. (3) The content should be of sufficient granularity for industrial practitioners, researchers and market surveillance authorities to derive required action or, in some cases, even reproduce the incident. The taxonomy should further be easy to use for reporters, and contain machine readable, properties that however depend on the exact implementation. Finally, we present this draft to spark further discussions in the community.
Outline. We first define relevant terminology based on existing standards or works, and review related concepts and approaches (Section 2). Afterwards, we describe how we derived our taxonomy by unifying related work and expert opinions with iterative rounds (Section 3). We present the resulting scheme in Section 4, and review limitations, implementation, implications and future work (Section 5).
Before concluding with the impact statement, we summarize alternative views (Section 6).
Related Work. To the best of our knowledge, there is no established procedure for AI security incident reporting. Cattell & Ghosh (2024) propose a CVE-like framework for adjudicating AI flaws more generally (e.g., incident reporting for undesired behaviors of AI systems which is not focused on security) as well as other improvements (such as extended model cards). In this sense, Cattell & Ghosh (2024) is orthogonal to our work. In addition, Fazelnia et al. (2024) proposed and MITRE announced a general AI incident sharing initiative [^4] https://ai-incidents.mitre.org concurrently to us. While both proposals bear similarities to ours, we suggest a more detailed treatment of the intended purpose, supply chains, and existing security measures. However, Fazelnia et al. (2024) focus on a minimum set of requirements and also propose an AI bill of materials (AIBOM). As AIBOMs summarize all components of AI models (Fazelnia et al., 2024; Bennet et al., 2024), our taxonomy would also benefit from an AIBOM standard.
2 Background
We first motivate our work based on the legal background. Afterwards, we introduce, based on existing standards, relevant definitions. We then review existing works describing AI systems, in AI security, before we discuss existing infrastructure to report AI incidents and security vulnerabilities.
2.1 Incident reporting as regulatory requirement
From a legal perspective, there are security-related requirements applicable under Article 55 (1c) of the European AI Act [^5] https://eur-lex.europa.eu/legal-content/EN/TXT/PDF/?uri=OJ:L_202401689. More specifically, commitment 17 on serious incident reporting of the general-purpose AI code of practice, currently under development [^6] https://digital-strategy.ec.europa.eu/en/library/second-draft-general-purpose-ai-code-practice-Dpublished-written-independent-experts, has a direct relation to our work: According to these documents, providers of general-purpose AI systems will be obliged to report relevant incident information such as affected models, root causes, or fail-safe measures in case serious incidents occur. With our taxonomy, we aim to cover security-related serious incidents. Our goal is to further to sharpen content, scope, and granularity of future reporting requirements. Especially as market surveillance authorities such as the European AI Office will have to issue further guidance on their operationalization, we want to initiate and inform a practical debate.
2.2 Definitions
Consequently, we review core terminology based on existing standards and use them to construct a definition of “AI security incident.” The complete list of all required definitions is in Appendix A. Here, we only revisit the three most important definitions: firstly an “AI system”, secondly “security”, and thirdly an “AI security incident” (the target of our collection). We first start with the definition of the “AI system”, derived from the OECD (a) definition.
Definition 2.1.
AI system. An AI system is a machine-based system that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments. Different AI systems vary in their levels of autonomy and adaptiveness after deployment.
We look to this definition of “AI system” as a foundation for our work for two main reasons. Firstly, the relevant OECD recommendation also identifies that “AI systems” and the “AI models” they embed are distinct. This allows a potential AI security incident report to identify if an exploit is scoped strictly to the “mathematical construct” that is a model, or interacts with the larger AI system. Secondly, the definition is used by or influences those of other intergovernmental organizations (such as the United Nations) and prominent market regulations relevant to incident reporting (such as the EU AI Act) (OECD, a).
We also recognize “security” as broadly defined in ISO/IEC 27000 by ISO (2018).
Definition 2.2.
Security. The preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of a system.
We further additionally review the definitions of confidentiality, integrity, and availability, also known as the “CIA triad”. This well established concept is also used in scientific work covering AI security (Biggio & Roli, 2018).
Definition 2.3.
Confidentiality. Property that information is not made available or disclosed to unauthorized individuals, entities, or processes (ISO, 2018).
Definition 2.4.
Integrity. Property of accuracy and completeness (ISO, 2018).
Definition 2.5.
Availability. Property of being accessible and usable on demand by an authorized entity (ISO, 2018).
Importantly, this definition of “security” based primarily on the CIA triad is shared by other ISO/IEC standards specific to AI. [^7] For example: ISO/IEC 42001 Information technology — Artificial intelligence — Management system; ISO/IEC TR 24028 Information technology — Artificial intelligence — Overview of trustworthiness in artificial intelligence; ISO/IEC TR 24368 Information technology — Artificial intelligence — Overview of ethical and societal concerns
Based on these foundational terms, we define an “AI security incident”, extending the existing definition of “AI incident” by the OECD (2024) and considering the above CIA security triad aspects shared across ISO/IEC standards.
Definition 2.6.
AI security incident. An AI security incident is an event, circumstance, or series of events where the confidentiality, integrity, or availability of an AI system is actually or potentially affected and which leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labor and intellectual property rights; (d) harm to property, communities or the environment.
We use these definitions to avoid, for example, confusion between general security and AI-specific security incidents (Bieringer et al., 2022). Also note that the proposed definition of an AI security incident includes potential compromises, hence vulnerabilities or red teaming results. It may be desirable to formulate this more strictly later-on.
2.3 Describing AI systems
Since the reporting of AI security incidents should be consistent with the overall documentation of an AI system, our proposed taxonomy builds on and expands prior work on model cards. After their initial proposal by Mitchell et al. (2019), these have been widely adopted by model providers and are used by various stakeholder groups to access information about a model and conclude corresponding action. Derivative templates have been proposed as fact sheets (Arnold et al. (2019)), datasheets for datasets (Gebru et al. (2021)), method cards (Adkins et al. (2022)), risk cards (Derczynski et al. (2023)), use case cards (Hupont et al. (2024)), and AI cards (Golpayegani et al. (2024)). While most of these approaches require the documentation of general metadata about the model (e.g., model type or model version), their content related to security is often limited and only covers implemented security measures, if at all. Yet, they provide suitable guidance on how to describe the underlying system affected by the incident. This is why the descriptions of AI systems affected by incidents within our proposed taxonomy (Section 4.2) are based on this body of literature.
2.4 AI security
Similar to any other (digital) system, an AI system can be targeted by a plethora of attacks (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). To understand how AI is vulnerable, in the following, we give a very high-level overview based on the previously discussed CIA triad defined in the previous Section 2.2.
Evasion attacks tamper with the test input to change the output of the model (Biggio & Roli, 2018; Dalvi et al., 2004; Boloor et al., 2020; Abdelnabi et al., 2023), thus harming the model’s integrity.
Poisoning attacks alter the training data, samples, or labels to decrease the overall performance of the classifier (Barreno et al., 2006; Cinà et al., 2023; Zhang et al., 2020). They, thus, target the integrity via the performance, or availability of the AI model if performance becomes too low.
Table 1: Existing databases and collections of AI (security) incidents. We list the kind of support party (Government, Industry, Non-Profit, and Academia) and incident numbers as of June 14th, 2024. We list the number of entries in each database and whether incidents cover the topics of ethics, security, AI security, or offensive use of AI within security. We then review which databases contain reproducible incidents. We then revisit the sources of the incidents: research papers, media, anonymous reports, or platforms where registered users can submit reports. We finally report whether the databases are machine-readable and linked with the CVE databases. The AVID database is linked to NVD, MITRE, and AIID.
Incident Type Sources
Database # Entries
Ethical
Security
AI Security
Offensive
AI usage
Reproducible
Incidents
Research
Paper
Media
Anonymous
Submissions
Machine
readable
Direct
Integration
Linked with/to
OECD (b) AIM G $>$1,000 CSV
Huntr I $>$1,000 CVE
AIID (McGregor, 2021) NP 691 CSV MongoDB
Deepfake DB (Walker et al., 2024) A 200 CSV
AVID (ARVA, ) NP 48 JSON Python
LVEs A 36 JSON Python
Grosse et al. (2024b) A 32
MITRE case studies (MITRE, a) G 22 JSON Python CVE
AIRD (MITRE, ) I 4 CSV Python CVE
Model stealing copies the model without the owner’s consent (Carlini et al., 2024b; Lowd & Meek, 2005; Oliynyk et al., 2023; Tramèr et al., 2016), targeting confidentiality.
Membership inference infers information about the training data based on the model’s replies to queries (Shokri et al., 2017; Pan et al., 2019; Fu et al., 2024). This attack thus also targets confidentiality.
More recent attacks include backdoors (Cinà et al., 2023; Gu et al., 2017; Yang et al., 2024; Yu et al., 2022), weight perturbations (Breier et al., 2018; Neti et al., 1992), adversarial initialization (Grosse et al., 2020; Liu et al., 2019), and sponge attacks (Shumailov et al., 2020). The exact nature and scope of describing and tagging a specific attack in a practical context is a research question itself and thus left for future work. There are, however, established frameworks like the MITRE ATLAS [^8] https://atlas.mitre.org/matrices/ATLAS addressing this challenge in practice. An alternative route is to use existing research frameworks, for example by Biggio & Roli (2018).
2.5 AI security incident reporting frameworks
Although there are reports of real-world AI security incidents as discussed in the introduction (Grosse et al., 2024b; McGregor, 2021; ARVA, ), there are not many security incident databases for AI systems. We summarize related databases in Table 1. Most databases cover not only AI security incidents but also other types of incidents, from unethical applications of AI to security; some include offensive use of AI (OECD, b; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a). To enable conclusions specific to AI security, databases should be more focused. This is the case for fewer databases, which are narrow but however collect non-AI security (Huntr, ) or offensive AI usage (Walker et al., 2024). To the best of our knowledge only Grosse et al. (2024b) focus on AI security, with only a small sample of 32 incidents. In addition, most databases rely on publicly available information, like research papers and media reports (OECD, b; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a). To enable the collection of AI security-relevant information, reports should be used that share relevant details. Yet, few databases rely on direct submissions (Huntr, ; LVEs, ; MITRE, ), including, for example, bug bounties. Even fewer use anonymous reports (Grosse et al., 2024b). Many collections also fall short in the underlying taxonomy that would capture relevant information for systematic analysis benefiting all stakeholders. While there are approaches with taxonomies, like the AI Incident Database (AIID) (McGregor, 2021), their focus should ideally be more narrow to allow security-specific conclusions. A suitable approach is the concurrently developed AI incident sharing initiative by MITRE [^9] https://ai-incidents.mitre.org. In contrast to this initiative, we advertise sharing mode details for analysis.
Still, many databases provide guidelines on collected information and emphasize other relevant features. These include machine readability (seven of nine databases allow direct export of data), links to other vulnerability databases like CVE (four of nine), and lastly direct integration via for example APIs with programming languages for data analysis (five of nine). The latter is especially relevant to research, where corresponding frameworks in non-AI security have enabled more detailed analysis of the threat surface (Neuhaus & Zimmermann, 2010; Li et al., 2017).
2.6 Security incident reporting frameworks
Frameworks for collecting, classifying, and reporting IT security incidents began to take shape in 1999 with the seminal work on the Common Vulnerability and Exposures (CVE) database (Baker et al., 1999). The CVE Program (CVE, ) has since matured and achieved widespread adoption. Building on the success of CVE, complementary initiatives evolved.
Common Vulnerability Scoring System (CVSS) (NIST, a; Scarfone & Mell, 2009) was developed to assess the severity of vulnerabilities. As a result, the National Vulnerability Database (NVD) (NIST, b) now contains CVSS scores for all published CVE records.
Common Weakness Enumeration (CWE) (MITRE, b) lists software and hardware weaknesses that could potentially evolve into vulnerabilities. There may be a connection between CVE and CWE records, as a reported CVE vulnerability could be the result of a documented CWE weakness.
MITRE ATT&CK (MITRE, b, c) is a knowledge base of attack tactics and techniques. It relates to CVE and CWE in that ATT&CK entries may exploit reported vulnerabilities and weaknesses.
Previous research has highlighted that security vulnerabilities in general and those specific to AI are fundamentally different (NIST, 2023; Cattell & Ghosh, 2024). We agree with this position, yet acknowledge that the CVE project can also be relevant for AI models. For example, CVE-2023-25661101010https://nvd.nist.gov/vuln/detail/CVE-2023-25661 reports a Tensorflow-related vulnerability that could have implications for AI models. However, the scope of CVE, and its associated taxonomy, differs from our approach. The CVE project primarily focuses on IT-related vulnerabilities, while our AI security incident taxonomy is specifically designed to address incidents related to AI models.
Our work aims to translate subject matter expert input into a comprehensive, self-contained AI security incident taxonomy. This approach mirrors the CVE model, where reported vulnerabilities are reviewed and eventually added to the CVE database. This also means that our project can benefit from lessons learned throughout the evolution of the CVE project. For instance, as the volume of AI security incidents is expected to grow steadily, a labeling schema is required to uniquely identify incidents that are classified according to our proposed taxonomy. Additionally, the process of reviewing and labeling incidents could become resource-intensive. Therefore, a governance body — similar to the CVE Numbering Authorities (CNA) (MITRE, a) — must be established to oversee this task.
If our proposed taxonomy gains acceptance within both the AI and security communities, further initiatives can be explored such as integrating the CVE and our taxonomy, assigning severity scores similar to CVSS, or abstracting AI security incident instances into more generalized weaknesses, similar to those defined by CWE.
3 Approach
To derive a taxonomy for reporting and describing AI security incidents, we followed a three-step approach. Firstly, we reached out to twelve industry experts (working in AI red teaming, security industry, AI governance, or market oversight), inquiring about their wish list for a possible reporting scheme. After distilling concepts from these conversations, we matched these concepts across experts, with model card approaches (Section 2.3), and existing databases (Section 2.5). The derived, rudimentary scheme was then refined by iterative discussion rounds among four of the authors. As a consequence, duplicates were removed. For example, some experts raised the question of whether the model was re-used, while others noted the relevance of keeping track of supply-chain-related information regarding the model. Here, the fourth-fold structure of the taxonomy emerged. In this process, we added meta-information from CSET (Hoffmann et al., ) like incident date, annotator, and whether the AI system was applied in a cyber-physical context.
When both, taxonomy draft and definitions, had sufficiently converged, three of the authors used the tentative structure to tag incidents from existing databases. The tagging was overall consistent, except in fields like the classification of harm, where two authors had significantly less background than the third; showing the need for guidelines. Other examples with disagreement related to the generalizeability of incidents, a challenging question for which, to the best of our knowledge, standard assessment procedures are not yet established. Afterwards, the proposal (in the form of this paper) was shared with the initial and additional experts for a broader, final round of feedback. Here, no significant changes to the taxonomy occurred.
4 AI security incident taxonomy
Our proposed taxonomy consists of five parts, where the last contains only an open text field for further comments and feedback. The parts are, (1) the incident’s meta data, (2) information about the targeted AI system, (3) specifics of the security incident, and lastly (4) implications of the incident. We now revisit the individual questions and references in each part. An example of the final questionnaire is depicted in Table 2-4 in the appendix and Figure 1 gives an abstract overview on our proposed taxonomy.
Figure 1: Abstract template of our proposed taxonomy for reporting and describing AI security incidents.
4.1 Meta data (T1.1-T1.4)
The metadata is a short block containing the incident’s high-level information, and is, under variations, part of many existing databases (Huntr, ; McGregor, 2021; LVEs, ; ARVA, ; OECD, b). This information forms the first section with taxonomy elements T1.1-T1.4.
We opted to assign an incident ID, a name or ID of the incident reporter, their organization, and the alleged date (or interval, or unknown) of the incident. It is important to remark that an AI security incident may be reported anonymously. Although, as stated above, any question of our proposal can be dropped, we want to emphasize that the reporter and their organization may be pseudonymized or kept fully anonymous, depending on the context of usage.
4.2 Targeted AI system
To understand an AI security incident, it is crucial to describe the AI system that formed part of the incident. Hence, in the second section of our taxonomy, we first collect information about the model, its life cycle stage and use, its origin, and application-specific information. Ideally, this information is directly imported from a model card or similar documentation to reduce required efforts.
Model information (T2.1-T2.3). We first collect basic information about the model. This includes the model type, for example, whether a model is generative or predictive; a specific reply here could be a support vector machine. Such information was suggested by half of our experts and is used in several model card approaches (Mitchell et al., 2019; Arnold et al., 2019; Adkins et al., 2022). We then collect more specific information like the exact version of the model, information also encompassed in many model cards (Mitchell et al., 2019; Adkins et al., 2022) and requested by a fourth of our experts. Lastly, if possible, the model should be specified as detailed as possible (e.g., model weights), to enhance reproducibility, as suggested by another quarter of our experts.
Life-cycle stage and intended use (T2.4-T2.7). In addition to the model within the AI system, understanding the system context is crucial. To this end, we also collect lifecycle information as specified by ISO [^11] ISO/IEC 22989 Information technology — Artificial intelligence — Artificial intelligence concepts and terminology, https://www.iso.org/standard/74296.html, as suggested by two of our experts. As suggested by the CSET AI Harm Framework (Hoffman & Frase, 2023) annotation guide, we also collect the domain where the AI system is deployed (Hoffmann et al., ). Intuitively, a deployment in physical domains has different implications than in the digital space. In addition, we keep track of the intended use of the AI system, loosely following the EU AI Act (Council of European Union, 2024), and as done in many model card approaches (Mitchell et al., 2019; Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). To take into account specific exploits, we also advocate denoting a usage during the incident if different from the intended use.
Model supply chain (T2.8-T2.10). Furthermore, it is crucial to understand possible (external) influences on the AI system stemming from the supply chain, as also suggested by two of our experts and in several AI model card approaches (Mitchell et al., 2019; Gebru et al., 2021; Adkins et al., 2022). Following scientific literature of AI security, we opted to divide the supply chain into three major areas: data (Mir, 2024), model weights (Gu et al., 2019), and code (Grosse et al., 2020; Shumailov et al., 2021). Any of these may allow the import of security vulnerabilities. Especially for data, prior work suggests which information could be collected (Bender & Friedman, 2018).
Application (T2.11). Instances of AI security incidents may need specific details to be reproducible or understandable. As an example, we suggest collecting the system prompt in the case where a large language model (LLM) is used. Such information was suggested by one expert, is collected in LLM-related databases (LVEs, ), and model cards (Derczynski et al., 2023).
4.3 Incident information
Having characterized the AI system, we are ready to collect information about the security incident with the goal of understanding what (may have) happened. To this end, we first collect general information about the incident, then report relevant affected components and finally understand which defensive measures were in place.
General information (T3.1-T3.4). First, the materialization status is collected, as suggested by two experts and as done in one database (ARVA, ). This information distinguishes a vulnerability found during testing (e.g., red-teaming) from an exploit of a deployed system. We further put forward attempting to find the root cause of the incident and report it. To ensure that an incident is indeed an AI security incident as defined in Definition 2.6, the attacker’s malicious intent should be recorded. This was suggested by half of our experts, forms part of several databases (ARVA, ; OECD, b; Walker et al., 2024), and was suggested in prior work (Grosse et al., 2024b). Finally, we recommend recording the attack type. For our tagging, we relied on the categorization by Grosse et al. (2024b). Any other framework, however, may also serve this purpose.
Relevant components (T3.5-T3.7). An AI security incident is further categorized by the resources used by the attacker to carry out the attack (Apruzzese et al., 2023). We thus recommend identifying and reporting the target asset, e.g. whether the model or the entire system was targeted, pointed out in prior work (Bieringer et al., 2022). In addition, we propose to keep track of used hardware, as suggested by one expert. Finally, access rights to all components of the AI system (data, model, queries, etc) should be reported to identify how the attacker possibly entered the system. The latter was suggested by one expert and studied in prior work (Grosse et al., 2024a, 2023).
Defensive measures (T3.8-T3.9). Access patterns also give an understanding which access may have prevented the attack. Along these lines, there is value in understanding what measures were in place before the incident took place, but did not prevent it. To this end, we advocate reporting existing security measures, as suggested by a quarter of our experts and the literature in model cards (Hupont et al., 2024; Arnold et al., 2019; Adkins et al., 2022). This is in particular relevant for large language models, where implemented guardrails should be reported (Rebedea et al., 2023). Along the same lines, we suggest also collecting fail-safe measures.
4.4 Incident implications
To understand the severity of an AI security incident, it is crucial to monitor resulting damage. We capture such effects in this brief section with elements T4.1-T4.4. The information in this section may be difficult to derive depending on the role of the reporter—a possible solution for this problem is to rely on a committee or internal risk management teams.
As the first element, we propose to monitor whether the incident led to non-compliance with existing regulations, a factor brought up by industry participants in security contexts (Bieringer et al., 2022; Grosse et al., 2023). Secondly, any harm that occurred should be well-documented, as suggested by the CSET AI Harm Taxonomy (Hoffman & Frase, 2023) applied within the AI Incident Database (McGregor, 2021). While determining (or defining) the extent of harm of an AI incident can be inherently uncertain (Paeth et al., 2024), we advocate that it is still worth attempting to record harms resulting from security incidents.
A different level of abstraction is to use the CIA definition from Section 2.2 to characterize the impact of the incident. Finally, and beyond the immediate incident, the goal of our taxonomy is to understand vulnerabilities beyond an individual system. We thus advocate explicitly reporting an assessment of whether the incident could affect other AI systems. This was also suggested a fourth of our experts and is implemented in the Huntr database already (Huntr, ).
4.5 Appendix
As a last part, we propose a field for links to connect the reported incident to related incidents from either a different database or a similar incident from the same reporting scheme, via either name or an URL. Finally, we suggest adding a comment field. This allows adding relevant information missing in the previous parts. Such a field avoids loss of information while seeking to keep the reporting taxonomy directed and usable.
5 Limitations, implementation, implications and future work
Having discussed our taxonomy, we revisit its limitations or possible extensions. We then discuss desirable features when implementing the taxonomy, for example, machine readability. We conclude with related open questions, for example the tension between intellectual property and the information collected.
Limitations. Our taxonomy is a basis for a discussion within the community. As such, it has several limitations. Firstly, it does not cover offensive AI, in other words using AI to carry out attacks. Such usage is, in our opinion, orthogonal to security incidents. Secondly, the proposal currently does not handle AI agents, multi agent systems, AI systems with multiple AI models or components, or recursive usage of AI. While an extension covering several AI systems is possible by filling questions $2.x$ for all covered AI systems, there may be better solutions to this shortcoming. In general, the taxonomy and underlying definitions will have to be adapted for new use cases and trends in AI. Furthermore, AI and its security incidents are very application specific [^12] As an example, consider AlphaFold’s prohibited use policy https://github.com/google-deepmind/alphafold3/blob/main/WEIGHTS_PROHIBITED_USE_POLICY.md, and it may be necessary to adapt the taxonomy for specific applications. Furthermore, our current definition of security also covers incidents with “potential” effects. While this enables to cover red-teaming results or vulnerabilities with unknown statuses of exploitation, in the long run it may be more beneficial to align the definition of security to include only incidents with identified harm.
Implementation. Many properties desired by the community, such as ease-of-use, machine readability (Section 2.5), and reproducibility (Section 4.2) depend on the practical implementation of our taxonomy. For example, the taxonomy is best to use when frameworks for attack and harm categorization are well known to the reporter and additional stakeholders involved in incident reporting and mitigation. We thus suggested that external and pre-existing frameworks in certain corresponding questions (for example, attacks and harm questions), can be utilized (Section 4.3). Along these lines, machine readability (Section 2.5) depends strongly on whether answers are selected from finite lists (such as multiple-choice or drop-down responses) as opposed to open text fields, or defining the taxonomy keys and values in a more generalizable schema (such as XML, RDFa, or other standards). While we most likely cannot eliminate all open text fields, we strongly suggest minimizing them as much as possible in order to allow ease of analysis. Lastly, it would be good to allow the reproducibility of an incident (Section 4.2). However, reproducibility directly contradicts the intellectual property of a model release. In this context, it would be possible to reduce the reporting scheme to contain only information supporting understanding of the incident, and not release model weights, for example. This would also include enforcing exact versioning to allow referencing a model without sharing it. Otherwise, a trusted third party could act as an intermediary.
Implications and future work. We have outlined that the exact reporting taxonomy and implementation strongly influence the overall usefulness and applicability of an AI incident reporting framework. Understanding the requirements of new emerging trends within AI (security), new applications, and, machine readability of AI incident data are open research questions. This extends to the tension between the protection of proprietary AI model information and the learning and generalizability of AI security incidents. The draft we presented here is intended as a scaffold to determine, discuss, explore and apply these facets.
Orthogonally, but analogous to other areas of incident reporting, usability studies could help to understand and advance both the usefulness of the collected reports (Marcilly et al., 2019; Mazaheri et al., 2015; Wiele & Rantanen, 2015). We have also outlined that in some cases, like supply chain (T2.8-T2.10, Section 4.2), attacks (T3.4, Section 4.3), and harm (T4.2, Section 4.4), a decision has to be taken which frameworks are to be chosen. This is an additional avenue for future work. Another aspect of our taxonomy could be to identify attacker’s identity. While similar frameworks such as CVE neither cover the identification or the collection of information on the attacker, in the long run approaches obtaining such information would greatly benefit both research, AI deployers and stakeholders.
6 Alternative views
The position and taxonomy proposed in this paper are not the only way to approach AI security incident reporting. While the exact taxonomy, as stated, is open for debate, there are other positions to be taken on this topic.
Position—No reporting is needed. One may argue that due to the absence of many real world AI security incidents, AI security reporting is not (yet) needed. In contrast, we believe that based on initial reports of AI security incidents (Grosse et al., 2024b; McGregor, 2021; ARVA, ) and the idea of proactive security (Biggio & Roli, 2018), appropriate incident collection and analysis is crucial. In other words, incident reporting should start now to enable proactive, as opposed not reactive security treatment.
Position—CVE reporting is sufficient. Another position is to integrate AI security reports into the CVE program(Fazelnia et al., 2024), possibly altering the CVE scheme to better fit AI security incidents. We have highlighted several counterarguments (Section 2.6), including that CVE is too general and does not consider that data and AI models are often withheld by companies. However, regulatory frameworks might require model providers and deployers to make the specific information related to AI security incidents transparent, enabling CVE to be used.
Position—Extending general AI incident reporting. A different approach to the one described above could be relying on existing infrastructure (Section 2.5) for non-security AI incidents and extending it to security. While we argue that the frameworks are sufficiently different and security incidents merit specific attention, existing infrastructure and AI incident communities could be extended to security.
Position—A smaller set of information is needed. There may be criticism that the over 25 questions (Section 4) are too many and a bureaucratic burden for the reporters. While we argue in favor of reusing existing infrastructure such as model cards (Section 4), we would like to emphasize our taxonomy is not a minimal, but an ideal suggestion to likely satisfy the requirements of all affected or involved stakeholder groups. We are open to discussions around and a creation of a minimal subset of our taxonomy.
Position–Taxonomies evolve over time. As reasoned above, security reporting infrastructures have been changing over time to accommodate emerging new methods and tools (Section 2.6). It is thus expected that a taxonomy of AI security incidents will evolve in a similar fashion over time. Yet, we argue that this draft is valid as a base of discussion. Eventually, a maintained framework is needed to standardize communication between reporting entities (e.g., providers of general purpose AI models with systemic risk) and entities receiving reports as required by applicable regulation (e.g., European AI Office).
Impact statement
The current annual damage of cyber crime, when quantified in monetary terms, was eight trillion USD in 2023 alone. [^13] https://www.statista.com/forecasts/1280009/cost-cybercrime-worldwide Understanding and combating such crime is thus of uttermost importance. While the majority of this number is currently caused by “classical” security issues (unrelated to AI), the deployment of AI systems across sectors is dramatically increasing and AI is likely to be implicated in far more cybercrime losses. We hope with our paper to make a significant step towards a taxonomy for AI security incidents that benefits all stakeholders—the developers and deployers affected, but also research, industry, and governmental entities who seek to analyze and learn from these incidents.
Acknowledgments We would like to thank our anonymous experts and their investment in giving initial inputs and feedback on this work. Some of our experts consented to being explicitly named, including In addition, the authors would like to thank Brian Pendleton, whom we thank especially for his feedback regarding CVE/CWE; Hyrum Anderson, Cisco; Leon Derczynski, NVIDIA Corporation, whom we thank for his feedback on our definitions, methodology, and references; and Yuvaraj Govindarajulu, AIShield, Bosch Global Software Technologies. In addition, we would like to thank Sean McGregor, AIID (AI incident database) for his feedback.
Appendix A Definitions
In the main paper, we list the most important definitions. To define all aspects of our taxonomy well, we introduce all required definitions and their sources in the following section.
Definition A.1.
Vulnerability. A weakness of an asset or a measure to reduce risk that can be exploited as a potential cause of an unwanted incident, which can result in harm to a system or organization, from ISO/IEC 27000:2018 (ISO, 2018).
Definition A.2.
Attack. Attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of an asset, from ISO/IEC 27000:2018 (ISO, 2018).
Definition A.3.
AI system security. Preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of an AI system, derived from security definition (Def. 2.2).
Definition A.4.
AI security. Preservation of confidentiality, integrity, availability, accountability, authenticity, and reliability of an AI model, derived from Security definition (Def. 2.2).
Definition A.5.
Incident. An incident is an event, circumstance or series of events where the development, use or malfunction of one or more systems directly or indirectly leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, from the OECD [^14] https://www.oecd-ilibrary.org/docserver/d1a8d965-en.pdf?expires=1718528495&id=id&accname=guest&checksum=8938FF35EAD65EAADD585DCF1FE65791.
Definition A.6.
AI Incident. An AI incident is an event, circumstance or series of events where the development, use or malfunction of one or more AI systems directly or indirectly leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, from the OECD [^15] https://www.oecd-ilibrary.org/docserver/d1a8d965-en.pdf?expires=1718528495&id=id&accname=guest&checksum=8938FF35EAD65EAADD585DCF1FE65791
Definition A.7.
Non-security incident. Any incident that was not caused by an attack and harmed the security of a system, based on definitions for security (Def. 2.2) and attack (Def. A.2).
Definition A.8.
Non AI-security security incident. A non AI security incident is an event, circumstance or series of events where the confidentiality, integrity, or availability of an AI system, but not the ML model, is actually or potentially affected and which leads to any of the following harms: (a) injury or harm to the health of a person or groups of people; (b) disruption of the management and operation of critical infrastructure; (c) violations of human rights or a breach of obligations under the applicable law intended to protect fundamental, labour and intellectual property rights; (d) harm to property, communities or the environment, derived from the definitions of AI security, security, and attack (Def. A.2).
Definition A.9.
Offensive use of AI. The use or abuse of ML to accomplish a malicious task, as defined by Mirsky et al. (2023).
To distinguish AI system and AI model, we rely on the reasoning of the OECD (a): “Although different interpretations of the word “model” exist, in this document, an AI model is a core component of an AI system used to make inferences from inputs to produce outputs.”, referring to the following technical definition:
Definition A.10.
AI model. Mathematical construct that generates an inference or prediction based on input data or information, from ISO/IEC 22989:2022.
Definition A.11.
Data. Re-interpretable representation of information in a formalized manner suitable for communication, interpretation or processing, from ISO/IEC TR 24028:2020.
Definition A.12.
Infrastructure. Hardware and software environment to support computer system and software design, development, and modification, from ISO/IEC TR 17903:2024.
Definition A.13.
Safety. Freedom from risk which is not tolerable, from ISO/IEC TR 5469:2024.
Table 2: Reporting scheme with description for annotators and examples, first third (Part 1.1 - Part 2.11).
Short name Long name or question Description for annotators Example
Meta Information T1.1 Incident ID Incident ID The unique identifier of the incident. ID_1716
T1.2 Incident Reporter Name of incident reporter Name or also username from within recording. John Doe / AIIncidentreporter001
T1.3 Reporting Organization Name of reporting organization The legal name of the reporting organization. Acme Inc.
T1.4 Incident Date Date the incident ocurred Date, Time interval, if unkown estimate interval 13/12/2024
AI System Information T2.1 Model type Specify the ML model type. Generative or predictive AI? prediction, generative, but also more concrete, Transformer, DNN, ConvNet, SVM etc Support Vector Machine
T2.2 Model version High level identifier, as noted in model card. This identifier should allow for correct identification of the specific model version. Llama3.2:1B
T2.3 Model specification Provide as precise information on the model as possible. Information could contain URLs to or files with weights, full model files or model cards. ’URL’
T2.4 Lifecycle stage Define the stage of the AI lifecycle which the model was in when the incident occurred. Stages of the AI lifecycle are: Inception, Design & Development, Verification & Validation, Deployment, Operation and Monitoring, Re-evaluate, Retirement. Deployment
T2.5 Domain Did the incident occurr in a physical environment? Did the incident affect an AI that is interacting with physical objects or otherwise physically with humans? Yes, pedestrian not detected.
T2.6 Intended use Described the intended use of the AI system. Intended purpose means the use for which an AI system is intended by the provider, including the specific context and conditions of use, as specified in the information supplied by the provider in the instructions for use, promotional or sales materials and statements, as well as in the technical documentation. Object detector to tag cats on image, Chatbot to reply to questions about air plane fares, etc
T2.7 Actual use during incident, if differing from 2.6 Describe the use of the AI system during the incident, if differeing from intended use (2.6) This may mean the usage for which an AI system was not intended by the provider, including the specific context and conditions of the actual use. Used to tag foxes instead of cats, used to get legal advice model was not trained for.
T2.8 Model supply chain - Data Was publicly available data used? If yes, from where, was data extended or adapted and if yes how? Specify dataset with name and URL, as well as any modifications that the data underwent before being used. For example: Model used MNIST data from ’URL’, where we added 1000 images that were handlabelled by our team.
T2.9 Model Supply Chain - Model weights Are the model weights derived from an existing model, and if yes, which one and how were they adapted? If the original design or weights have a name, provide this name or possibly a URL. Further describe all steps the model undewerwent to adapt it to the current status. For example: Based on Llama3.2:1B, but finetuned on the above described dataset for n iterations using optimzier o and loss l.
T2.10 Model Supply Chain - Code What public code was used to generate and run the model? Provide libraries used with their version name, and also ennumnerate github repositories with link and access data or other re-used code from the project. For example: Pytorch version 2.4.1. and Code from github repo ’this url’ was used, forked January 1st 2024.
T2.11 System prompt If the system is a generative AI system, provide the system prompt. If known, report the entire system, prompt as used in the setting where the incident occurred. For example: ’please be polite to the user and don’t hallucinate. The user writes: ’
Table 3: Reporting scheme with description for annotators and examples, second third (T3.1)-T3.9.
Short name Long name or question Description for annotators Example
Incident specification T3.1 Materialization status Has the incident been reported prior or after its materialization? Was the attack found during red-teaming or scientific analysis, or as part of an exploited breach? Red teaming
T3.2 Root cause Describe the root cause of the incident. What is the underlying reason, design choice, or property of the system that enables the attack? ’Attacker relied on infinity norm attack the model wasn’t tested for’, or ’API allowed infinite amount of queries, hence enabling crating a copycat training dataset’
T3.3 Malicious Intent Is there evidence that an entity or a person purposely caused the incident by targeting the AI systems security? Is the incident an example of an attempt to destroy, expose, alter, disable, steal or gain unauthorized access to or make unauthorized use of the AI system? ’Yes, the attacker altered the sample until it passed the AI security check’ or ’No, the user did not know that the system was not trained on this specific subset of the training data (Disease B). Hence, the system provided a wrong output.’
T3.4 Attack type Specify the concrete type of attack that has been carried out. Any attack framework, tagging based on the attacks in Sect. 2.4 Poisoning
T3.5 Target asset Did the attack target the model explicitly and the model only? Based on the definition in Sect. 2.2 Yes, model was targeted.
T3.6 Hardware What hardware was used, and could it have affected the incident? Please describe the specific hardware used on the device or server the application ran, and whether the setting implies that the hardware played a role in the incident. Mac M1, could not have affected the incident OR could have affected the incident via a glitch.
T3.7 Access Which access was given to whom for model, inputs, training, test or evaluation data? Specify which levels of access were given to which users regarding each component. Training and evaluation data as well as model was only available to internal model team. The model could be queried via API by the client, but access control prevented any other access by 3rd parties to the model.
T3.8 Security measures What security measures where used? Where they circumvented and if yes how? Give a list of specific security measures that were used including their configurations. Adversarial training using FGSM with epsilon 0.1, 0.2 and 0.3 was used on randomly selected 10% of the dataset were used. Defense was not circumvented as attack was based on L2 norm, not $L_{\infty}$ norm.
T3.9 Fail safe measures Where there any fail-safe measures implemented? Describe any fail-safe measures, such as rebooting the system, not outputting a classification, etc. System fell back on human intervention to prevent further harm.
Table 4: Reporting scheme with description for annotators and examples, last third (T4.1)-T5.2.
Short name Long name or question Description for annotators Example
Incident Implications T4.1 Regulatory violation List specific regulatory requirements that are violated due to the incident. Please specify any concrete legislation, standard or policy that has been violated by the incidents’ occurrence. European Union AI Act
T4.2 Resulting harm What harm did the incident cause to any participating stakeholder? Did the incident lead to any of the following types of harm? Affecting physical health, inducing financial loss, damaging a physical object, affecting intangible property, affecting infrastructure, affecting natural environment, violation of human rights, civil liberties, civil rights or democratic norms. Harm to physical property of all users, mainly food but also some devices due to malfunction of the fridge AI.
T4.3 Security Impact Describe how the incident affected the confidentiality, integrity, availability of the AI system. Describe whether the incident caused any leakage of confidential data, lead the system to output a wrong result or to exhibit a bias (integrity), or lead to the system being down or otherwise unusable (availability). During the incident, samples were manipulated to be classified as vehicles and thus harmed integrity.
T4.4 Generalization Does the incident potentially affect other AI systems? Incident descriptions should allow to determine for a third model if a vulnerability in that model is likely. Other AI systems might also be affected if relying on the same dataset ”xyz”.
Appendix T5.1 Links to other incidents Links to similar / same cases in other /same database(s) In case you are aware of a incidents that has been reported in another database or that is similar to this incident in this database This incident has had similar consequences like AIID incidents 38 or simply ’URL’
T5.2 Comments Additional comments. Anything you’d like to add which has not been covered in the previous fields. The attack does not work on model ABC1.2 and ABC1.4, but only on ABC1.3